# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 4.1. Deterministic approximation step: interpolation
The first (deterministic) approximation mimicks the approach of [KohnPhil24]. However, this approximation will not be useful in practice as it involves expectation values and thus is not computable, in general. Therefore, we merely use it as a theoretical tool for an error bound of a computable estimator which we shall propose in the next subsection.
@@
where $S_{\mathcal{X}}$ is the $\mathbb{H}$-orthogonal projection onto
$V_{\mathcal{X}}=\operatorname{span}\{\Phi_{x_{i}}:i\in\{1,\dots,d\}\}\subset\mathbb{H}$
with the canonical features $\Phi_{x_{i}}=k(x_{i},\cdot)\in\mathbb{H}$. With the notation from Section 2.3, recall the explicit projection formula (10) given by $S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ for some $g\in\mathbb{H}$.
Hence, application of the abstract operator defined in (31) may explicitly be computed via
$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=(\mathcal{K}S_{\mathcal{X}}f)_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\sum_{i=1}^{d}\big[\mathbf{K}_{\mathcal{X}}^{-1}\big(\underbrace{\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=\cdot]}_{\mathcal{K}S_{\mathcal{X}}f}\big)_{\mathcal{X}}\big]_{i}\Phi_{x_{i}}.$
Using linearity of the expected value and the projection formula (10), we get, for each $j\in\{1,\dots,d\}$,
$\mathbb{E}[S_{\mathcal{X}}f(Y)|X=x_{j}]=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}]=\mathbb{E}[\mathbf{k}_{\mathcal{X}}^{\top}(Y)|X=x_{j}]\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}}.$
Hence, defining the coefficients
$\alpha_{\widehat{\mathcal{K}}f}=\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}},\qquad\mathrm{where}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{i,j}=\mathbb{E}(k(Y,x_{i})|X=x_{j}),$
@@
where $\mathcal{B}_{\Omega}(x)=\{\Delta\in\mathcal{B}_{\Omega}:\rho_{x}(\partial\Delta)=0\}$ for $x\in\Omega$, where $\partial\Delta$ denotes the boundary of the Borel-set $\Delta$. Note that this holds in the deterministic case where $\rho_{x}=\delta_{F(x)}$ with continuous $F:\Omega\to\Omega$: Then, $\rho_{x_{0}}(\partial\Delta)=\delta_{F(x_{0})}(\partial\Delta)=0$ is equivalent to $F(x_{0})\notin\partial\Delta$, so $F(x_{0})$ lies either in the interior or the exterior of $\Delta$, and continuity of $F$ gives the convergence. By the Portmanteau theorem [Klen2008, Theorem 13.16], the condition (33) is equivalent to the fact that $\mathcal{K}f$ is continuous for each $f\in C_{b}(\Omega)$. Hence, if (33) is satisfied, the Koopman operator is a contraction from $C_{b}(\Omega)$ into itself as for $f\in C_{b}(\Omega)$,
@@
Under Assumption 4.1, i.e. $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b},C_{b})$, the approximant $\widehat{\mathcal{K}}$ from (32) admits the error bound
@@
The proof is a particular case of the proof of [KohnPhil24, Theorem 3.4] and we repeat it here for the sake of clarity. As $\mathcal{K}-\widehat{\mathcal{K}}=\mathcal{K}-S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}=\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})+(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}$, we get
@@
$\leq\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}).$
Under the assumption $\mathcal{K}\in L(C_{b},C_{b})$, $\mathcal{K}$ is a contraction, and for constant $f\equiv c\in\Omega$
@@
By Proposition 2.4, the invariance assumption $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the operator norm $\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}$ is bounded. Thus, it is in order to focus on the projection error $\mathcal{I}-S_{\mathcal{X}}$ in dependence of the point set $\mathcal{X}$. The projection error in (34) attains a particularly intuitive form when considering Sobolev spaces as RKHS. As stated before, Wendland and Matérn kernels both generate such Sobolev spaces, allowing for the application of results of approximation theory ([Wend04, Theorem 11.13]) to further specify Theorem 4.2 in these cases. Under the appropriate sufficient condition from Section 3, the order of convergence can be quantified in terms of the smoothness of the underlying kernels. We exemplarily state the result for Wendland kernels below.
@@
Let $l\in\mathbb{N}$ and define $\sigma_{n,l}=\frac{n+1}{2}+l$. Consider the Wendland kernel $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ that generates the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ as reproducing kernel Hilbert space with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the classical Sobolev norm $\|\cdot\|_{H^{\sigma_{n,l}}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (23) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\sigma_{n,l}$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (31) satisfies
@@
Under (23), the Sobolev space $H^{\sigma_{n,l}}$ is invariant under the Koopman operator $\mathcal{K}$. By [Wend04, Theorem 11.17] or the restatement [KohnPhil24, Theorem 5.1], we may bound the projection error by
@@
For a large data set $\mathcal{X}$, the kernel matrix $K_{\mathcal{X}}$ is typically badly conditioned, leading to numerical instabilities. To alleviate this, one might regularize the kernel-based interpolation problem as in [BoldPhil24] to get $R_{\mathcal{X}}^{\lambda}f:={\arg\min}_{{g\in V_{\mathcal{X}}}}\sum_{i=1}^{d}\left|f(x_{i})-g(x_{i})\right|^{2}+\lambda\|g\|_{\mathbb{H}}^{2}$ with regularization parameter $\lambda\geq 0$ and solution operator $R_{\mathcal{X}}^{\lambda}$. The linear solution operator satisfies $R_{\mathcal{X}}^{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$} for $f\in\mathbb{H}$, so for $\lambda=0$ we recover $S_{\mathcal{X}}$, but for $\lambda>0$ $R_{\mathcal{X}}^{\lambda}:\mathbb{H}\to\mathbb{H}$ is no longer a projection. However, as is shown in [BoldPhil24, Appendix A], in analogy to (31) we can still define the approximant
@@
Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}$ generated by the Wendland kernel $k_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, the deterministic error admits the bound
@@
The proof of (38) follows the lines of the proof of Theorem 4.2 and uses the fact that $\bigl\|R_{\mathcal{X}}^{\lambda}\bigr\|_{\mathbb{H}\to\mathbb{H}}\leq 1$ as shown in [BoldPhil24, Proposition B.1 (iii)]. The special case covered in (39) follows from [BoldPhil24, Theorem 2.4]. ∎
@@ 4.2. Probabilistic approximation step: estimating the expected value
In practice, we can not form $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$, as we do not have access to the expected values. For this reason, we introduce a random matrix ${\mathbf{K}}^{\mathrm{MC}}_{\mathrm{prop}}$, whose entries are Monte Carlo approximations of the expected values in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. We denote the $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ by $(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{j}$. The $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is then defined by
$\big(\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}\big)_{j}:=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\mathbf{k}_{\mathcal{X}}(y_{l,j})\approx\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)|X=x_{j}],$ (40)
@@
This leads to a computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ which is analogously defined as $\widehat{\mathcal{K}}$ in (32) with the only difference that we replace $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ with $\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$ as defined in (40), that is
@@
Decomposing the full error by inserting the auxiliary approximation $\widehat{\mathcal{K}}$, i.e,
@@
we are left with bounding the second term, after analyzing the deterministic first term in Section 4.1. Due to the random sampling employed in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$, this error is probabilistic. In the remainder of this part, we derive a probabilistic bound on the error
$(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}(\mathbf{K}_{\mathcal{X},\mathcal{Y}}-\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
@@
Let $\varepsilon>0$ and choose $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j=1,\dots,d$. Then, for all $f\in\mathbb{H}$, the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ from (41) satisfies
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\|f\|_{\mathbb{H}}$ (42)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$.
@@
$|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)|=|(g-g^{\mathrm{MC}})^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)|\leq\|g-g^{\mathrm{MC}}\|_{\infty}\,\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}.$
As shown in the proof of [BoldPhil24, Theorem 4.2], the deterministic second factor is bounded by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$. Moreover,
$\|g-g^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|.$
Setting $\mathbb{E}_{j}:=g_{j}=\mathbb{E}[(S_{\mathcal{X}}f)(Y)|X=x_{j}]$, we obtain $\mathbb{E}[(S_{\mathcal{X}}f)(y_{l,j})]=\int(S_{\mathcal{X}}f)(y)\,\rho(x_{j},dy)=(\mathcal{K}S_{\mathcal{X}}f)(x_{j})=\mathbb{E}_{j}$. Moreover,
$|(S_{\mathcal{X}}f)(y)|=|\langle S_{\mathcal{X}}f,\Phi_{y}\rangle_{\mathbb{H}}|\,\leq\,\|S_{\mathcal{X}}f\|_{H}\|\Phi_{y}\|_{\mathbb{H}}\,\leq\,\|k\|_{\infty}^{1/2}\|f\|_{\mathbb{H}}$
for all $y\in\Omega$, where the last inequality follows from $\|S_{\mathcal{X}}\|_{H\to H}=1$ since $S_{\mathcal{X}}$ is a projection. Hence, the random variables $(S_{\mathcal{X}}f)(y_{l,j})$ are bounded and i.i.d., so that we may apply Hoeffding’s inequality:
@@
$\leq 2\exp\left(-\frac{2m_{j}^{2}\varepsilon^{2}\|f\|_{H}^{2}}{4m_{j}\|k\|_{\infty}\|f\|_{H}^{2}}\right)=2\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
@@
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
@@
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H})$ $\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right).$
In total, for every $f\in H$, with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|k\|_{\infty}}\right)$, we obtain
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|k\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}\|f\|_{H},$
i.e., (42). ∎
Note that in the bound (42) we have $\|k\|_{\infty}<\infty$ by Assumption 4.1, since the kernel $k$ is continuous on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and hence bounded on the bounded open set $\Omega\times\Omega$.

### v2

@@ 4.1. Deterministic approximation step: interpolation
The first (deterministic) approximation mimicks the approach of [24]. However, this approximation will not be useful in practice as it involves expectation values and thus is not computable, in general. Therefore, we merely use it as a theoretical tool for an error bound of a computable estimator which we shall propose in the next subsection.
@@
where $S_{\mathcal{X}}$ is the $\mathbb{H}$-orthogonal projection onto $V_{\mathcal{X}}=\operatorname{span}\{\Phi_{x_{i}}:i\in\{1,\dots,d\}\}\subset\mathbb{H}$ with the canonical features $\Phi_{x_{i}}=\text{\sf{k}}(x_{i},\cdot)\in\mathbb{H}$. With the notation from Section 2.3, recall the explicit projection formula (11) given by $S_{\mathcal{X}}g=g_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}$ for some $g\in\mathbb{H}$. The action of the abstract operator defined in (32) now may explicitly be computed via
$\widehat{\mathcal{K}}f=S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}f=(\mathcal{K}S_{\mathcal{X}}f)_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}=\sum_{i=1}^{d}\big[\mathbf{K}_{\mathcal{X}}^{-1}\big(\underbrace{\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=\cdot]}_{\mathcal{K}S_{\mathcal{X}}f}\big)_{\mathcal{X}}\big]_{i}\Phi_{x_{i}}.$
Using linearity of the expected value and the projection formula (11), we get
$\mathbb{E}[S_{\mathcal{X}}f(Y)\mid X=x_{j}]=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)\mid X=x_{j}]=\mathbb{E}[\mathbf{k}_{\mathcal{X}}^{\top}(Y)\mid X=x_{j}]\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}}$
for each $j\in\{1,\dots,d\}$. Hence, defining the coefficients
$\alpha_{\widehat{\mathcal{K}}f}=\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}f_{\mathcal{X}},\qquad\mathrm{where}\qquad(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{i,j}=\mathbb{E}[\text{\sf{k}}(Y,x_{i})\mid X=x_{j}],$
@@
where $\mathcal{B}_{\Omega}(x)=\{\Delta\in\mathcal{B}_{\Omega}:\rho_{x}(\partial\Delta)=0\}$ for $x\in\Omega$, with $\partial\Delta$ denoting the boundary of the Borel-set $\Delta$. Note that this holds in the deterministic case where $\rho_{x}=\delta_{F(x)}$ with continuous $F:\Omega\to\Omega$: Then, $\rho_{x_{0}}(\partial\Delta)=\delta_{F(x_{0})}(\partial\Delta)=0$ is equivalent to $F(x_{0})\notin\partial\Delta$, so $F(x_{0})$ lies either in the interior or the exterior of $\Delta$, and continuity of $F$ gives the convergence. By the Portmanteau theorem [20, Theorem 13.16], the condition (34) is equivalent to the fact that $\mathcal{K}f$ is continuous for each $f\in C_{b}(\Omega)$. Hence, if (34) is satisfied, the Koopman operator is a contraction from $C_{b}(\Omega)$ into itself: for $f\in C_{b}(\Omega)$, we have
@@
Under Assumption 4.1, i.e. $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ and $\mathcal{K}\in L(C_{b}(\Omega),C_{b}(\Omega))$, the approximant $\widehat{\mathcal{K}}$ from (33) admits the error bound
@@
The proof is a particular case of the proof of [24, Theorem 3.4] and we repeat it here for the sake of clarity. As $\mathcal{K}-\widehat{\mathcal{K}}=\mathcal{K}-S_{\mathcal{X}}\mathcal{K}S_{\mathcal{X}}=\mathcal{K}(\mathcal{I}-S_{\mathcal{X}})+(\mathcal{I}-S_{\mathcal{X}})\mathcal{K}S_{\mathcal{X}}$, we get
@@
$\leq\|\mathcal{I}-S_{\mathcal{X}}\|_{\mathbb{H}\to C_{b}}\,\left(\|\mathcal{K}\|_{C_{b}\to C_{b}}+\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}\|S_{\mathcal{X}}\|_{\mathbb{H}\to\mathbb{H}}\right).$
Under the assumption $\mathcal{K}\in L(C_{b},C_{b})$, $\mathcal{K}$ is a contraction. Any constant $f\equiv c\in\Omega$ satisfies
@@
By Proposition 2.4, the invariance assumption $\mathcal{K}\in L(\mathbb{H},\mathbb{H})$ implies that the operator norm $\|\mathcal{K}\|_{\mathbb{H}\to\mathbb{H}}$ is bounded. Thus, it is in order to focus on the projection error $\mathcal{I}-S_{\mathcal{X}}$ in dependence of the point set $\mathcal{X}$. The projection error in (35) attains a particularly intuitive form when considering Sobolev spaces as RKHS. As stated before, Wendland and Matérn kernels both generate such Sobolev spaces, allowing for the application of results of approximation theory ([50, Theorem 11.13]) to further specify Theorem 4.2 in these cases. Under the appropriate sufficient condition from Section 3, the order of convergence can be quantified in terms of the smoothness of the underlying kernels. We exemplarily state the result for Wendland kernels below.
@@
Let $l\in\mathbb{N}$ and define $\sigma_{n,l}=\frac{n+1}{2}+l$. Consider the Wendland kernel $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ that generates the Sobolev space $H^{\sigma_{n,l}}(\Omega)$ as reproducing kernel Hilbert space $\mathbb{H}$ with equivalence between the native norm $\|\cdot\|_{\mathbb{H}}$ and the Sobolev norm $\|\cdot\|_{H^{\sigma_{n,l}}}$. Assume that $\mathcal{K}\in L(C_{b},C_{b})$ and that (24) is satisfied for some $\ell\in\mathbb{N}$ with $\ell>\sigma_{n,l}$. Then, the Koopman approximation $\widehat{\mathcal{K}}$ from (32) satisfies
@@
Under (24), the Sobolev space $H^{\sigma_{n,l}}$ is invariant under the Koopman operator $\mathcal{K}$. By [50, Theorem 11.17] or the restatement [24, Theorem 5.1], we may bound the projection error by
@@
For a large data set $\mathcal{X}$, the kernel matrix $K_{\mathcal{X}}$ is typically badly conditioned, leading to numerical instabilities. To alleviate this, one might regularize the kernel-based interpolation problem as in [7] to get $R_{\mathcal{X}}^{\lambda}f:={\arg\min}_{{g\in V_{\mathcal{X}}}}\sum_{i=1}^{d}\left|f(x_{i})-g(x_{i})\right|^{2}+\lambda\|g\|_{\mathbb{H}}^{2}$ with regularization parameter $\lambda\geq 0$ and solution operator $R_{\mathcal{X}}^{\lambda}$. The linear solution operator satisfies $R_{\mathcal{X}}^{\lambda}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}$} for $f\in\mathbb{H}$, so for $\lambda=0$ we recover $S_{\mathcal{X}}$, while for $\lambda>0$, $R_{\mathcal{X}}^{\lambda}:\mathbb{H}\to\mathbb{H}$ is no longer a projection. However, as is shown in [7, Appendix A], in analogy to (32) we can still define the approximant
@@
Specifically, for the RKHS $\mathbb{H}=H^{\sigma_{n,l}}(\Omega)$ generated by the Wendland kernel $\text{\sf{k}}_{n,l}$, where $\sigma_{n,l}=\frac{n+1}{2}+l$, the deterministic error admits the bound
@@
The proof of (39) follows the lines of the proof of Theorem 4.2 and uses the fact that $\bigl\|R_{\mathcal{X}}^{\lambda}\bigr\|_{\mathbb{H}\to\mathbb{H}}\leq 1$ as shown in [7, Proposition B.1 (iii)]. The special case covered in (40) follows from [7, Theorem 2.4]. ∎
@@ 4.2. Probabilistic approximation step: estimating the expected value
In practice, we cannot form $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$, as we do not have access to the expected values. For this reason, we introduce a random matrix ${\mathbf{K}}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$, whose entries are Monte Carlo approximations of the expected values in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. We denote the $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ by $(\mathbf{K}_{\mathcal{X},\mathcal{Y}})_{j}$. The $j$-th column of $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ is then defined by
$\big(\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}\big)_{j}:=\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}\mathbf{k}_{\mathcal{X}}(y_{l,j})\approx\mathbb{E}[\mathbf{k}_{\mathcal{X}}(Y)\mid X=x_{j}],$ (41)
@@
This leads to a computable approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$, which is defined analogously to $\widehat{\mathcal{K}}$ in (33), with the only difference that we replace $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$ with $\mathbf{K}^{\mathrm{MC}}_{\mathcal{X},\mathcal{Y}}$ as defined in (41). Thus, we set
@@
Decomposing the full error by inserting the auxiliary approximation $\widehat{\mathcal{K}}$, i.e.,
@@
we are left with bounding the second term, after analyzing the deterministic first term in Section 4.1. Due to the random sampling employed in $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$, this error is probabilistic. In the remainder of this part, we derive a probabilistic uniform bound on the error
$(\widehat{\mathcal{K}}-\widehat{\mathcal{K}}^{\mathrm{MC}})f=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\left(\mathbf{K}_{\mathcal{X},\mathcal{Y}}-\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\right)\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}.$
@@
Choose $m_{j}=m\in\mathbb{N}$ i.i.d. samples from every $\rho_{x_{j}}$, $j=1,\dots,d$, and let $\varepsilon>0$ such that $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$. Then, for all $f\in\mathbb{H}$, the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ from (42) satisfies
$\left\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\right\|_{\infty}\,\leq\,\varepsilon\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\sqrt{\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v}\cdot\|f\|_{\mathbb{H}}$ (43)
with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, where $\mathbbm{1}=\{v\in\mathbb{R}^{d}:|v_{i}|=1\}$.
@@
$\left|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f(x)\right|=\left|(g-g^{\mathrm{MC}})^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\right|\leq\left\|g-g^{\mathrm{MC}}\right\|_{\infty}\,\left\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\right\|_{1}.$
As shown in the proof of [7, Theorem 4.2], the deterministic second factor is bounded by $\|\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}(x)\|_{1}\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\big)^{1/2}$. Moreover,
$\|g-g^{\mathrm{MC}}\|_{\infty}=\max_{x_{j}\in\mathcal{X}}\left|\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=x_{j}]-\frac{1}{m_{j}}\sum_{l=1}^{m_{j}}(S_{\mathcal{X}}f)(y_{l,j})\right|.$
Setting $\mathbb{E}_{j}:=g_{j}=\mathbb{E}[(S_{\mathcal{X}}f)(Y)\mid X=x_{j}]$, we obtain $\mathbb{E}[(S_{\mathcal{X}}f)(y_{l,j})]=\int(S_{\mathcal{X}}f)(y)\,\rho(x_{j},dy)=(\mathcal{K}S_{\mathcal{X}}f)(x_{j})=\mathbb{E}_{j}$. We also have
$|(S_{\mathcal{X}}f)(y)|=|\langle S_{\mathcal{X}}f,\Phi_{y}\rangle_{\mathbb{H}}|\,\leq\,\|S_{\mathcal{X}}f\|_{H}\|\Phi_{y}\|_{\mathbb{H}}\,\leq\,\|\text{\sf{k}}\|_{\infty}^{1/2}\|f\|_{\mathbb{H}}$
for all $y\in\Omega$, where the last inequality follows from $\|S_{\mathcal{X}}\|_{H\to H}=1$ since $S_{\mathcal{X}}$ is an orthogonal projection. Hence, the random variables $(S_{\mathcal{X}}f)(y_{l,j})$ are bounded and i.i.d., so that we may apply Hoeffding’s inequality:
@@
$\leq 2\exp\left(-\frac{2m_{j}^{2}\varepsilon^{2}\|f\|_{H}^{2}}{4m_{j}\|\text{\sf{k}}\|_{\infty}\|f\|_{H}^{2}}\right)=2\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
@@
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
@@
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{\infty}\geq\varepsilon\|f\|_{H}\big)$ $\leq 2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right).$
In total, for every $f\in H$, we obtain the upper bound
$\|(\widehat{\mathcal{K}}^{\mathrm{MC}}-\widehat{\mathcal{K}})f\|_{\infty}\,\leq\,\varepsilon\cdot\|\text{\sf{k}}\|_{\infty}^{1/2}\cdot\Big(\max_{v\in\mathbbm{1}}v^{\top}\mathbf{K}_{\mathcal{X}}^{-1}v\Big)^{1/2}\|f\|_{H},$
i.e., (43), with probability at least $1-2d\exp\left(-\frac{m\varepsilon^{2}}{2\|\text{\sf{k}}\|_{\infty}}\right)$, which is nonnegative due to the imposed inequality $m\geq\frac{2\|\text{\sf{k}}\|_{\infty}}{\varepsilon^{2}}\log(2d)$. ∎
Note that in the bound (43) we have $\|\text{\sf{k}}\|_{\infty}<\infty$ by Assumption 4.1, since the kernel k is continuous on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ and hence bounded on the bounded open set $\Omega\times\Omega$.